OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 5) and (1, 4)? (1 point) –3 –2 –one third one third

OpenStudy (anonymous):

m=(5-4)/(-2-1)=-1/3

OpenStudy (anonymous):

(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

-ONE THIRD